Capacitors: 6 microFarad, 2 microFarad, and 6 microFarad combined in series give? ___ microFarad

Combined in parallel give? ___microFarad
Answer #1

Given q=6UF ; C2 = 2uF ; C3=64f combined in series 2 2 2 2 2 2 2 2 2 2 2 2 2 2 1 = 7 2 + 1 2 + 1/5 * = 1+3+ 1 = 56 of Cey= 6